### 22 Feb 2024 v2

1 Introduction
The estimation of quantile treatment effects in the presence of endogeneity has received increasing attention.
For this purpose, the instrumental variable quantile regression (IVQR) model introduced by [7] has been widely used in applications.
While the IVQR model assumes that outcome variables are one-dimensional, researchers are often interested in multiple variables at the same time.
For example, there is much work that studies generalized Engel curves to investigate how the shares of food and leisure expenditures change as total expenditures change. (See, for example, [23], [2], and [22]).
In such cases, researchers typically run the IVQR separately for each dimension of the outcome variables.
This approach is often problematic because it ignores how different variables are correlated with each other.
In the Engel curve example, the shares of food and leisure are significantly correlated because there is an obvious substitutability, i.e., more food, less leisure, so that they should not be treated independently.
In order to address this problem, we propose a novel quantile model that allows for multidimensional outcome variables that are potentially correlated with each other.
Consider the potential outcome framework of [7].
For a treatment let be the corresponding -dimensional potential outcome.
We assume that conditional on covariates, it is represented as where is a structural function and is a -dimensional rank variable vector.
Existing work on endogenous quantile models imposes different monotonicity assumptions for identification.
We generalize the assumption of [7] and assume that the structural function is cyclically monotone.
Under this restriction, an increase in the th entry of leads to an increase in the th entry of which justifies calling the rank variable vector of
More importantly, a change in a particular entry of can affect other dimensions of both positively and negatively.
This allows the model to capture substitutability/complementarity between different dimensions.
Running the standard IVQR separately does not have this flexibility.
A takeaway from [7] is that in order to identify quantiles of one-dimensional potential outcomes for binary treatments, an IV with binary support that satisfies the full rank condition is necessary.
For -dimensional potential outcomes, it may seem natural that the support size of an IV needed for the identification is greater than two because there are functions to be identified.
Surprisingly, however, our main theorem states that an IV with only binary support may achieve the desired identification.
At the cost of multidimensionality, the IV is required to satisfy what we call the positive definiteness condition, which is stronger than the full rank condition, and it becomes more demanding as the dimension increases.
The structural function of our model turns out to satisfy a system of functional equations.
We exploit this fact to show the identification.
Instead of showing the identification in the IVQR model directly, we give a general result that applies to any identification problem of potentially infinite-dimensional parameters that are characterized by a system of equations, which is of independent interest.
Related Literature:
The model in this paper stands on the IVQR model that is proposed and developed by [7], [8] and [10].
There are different studies that deal with endogenous quantile models when outcome variables are one-dimensional.
[1] consider quantile treatment effects on compliers, and [22] take the control function approach under the assumption that the selection equation is monotone.
A little work addresses potentially multidimensional outcome variables.
[3] propose a nonseparable model with multidimensional unobservable heterogeneity when a proxy variable of the heterogeneity is available.
[19] extends the control function approach to the multidimensional outcome setting.
Technically, the cyclical monotonicity assumption on the structural function requires it to be a multivatriate quantile function that is defined and developed by [6], [20], and [17].
Multivariate quantile functions are closely related to the optimal transport theory ([29] and [30]) and a partial differential equation (PDE), called the Monge-Ampère equation ([16]).
Indeed, the structural function of our model turns out to be a solution to a system of Monge-Ampère equations.
The local identification of the structural function can be interpreted as the local uniqueness of the solution to the system, which is one of the central issues in the PDE literature.
We prove the local identification result exploiting the positive definiteness of a differential operator, which is a powerful technique used in the elliptic PDE literature ([18]).
Theorem 5.1 in Section 5 generalizes nonparametric local identification results of [5] in that ours can be applied to any parameters that are characterized as the solution of a system of equations and that it does not require the parameter space to be a Banach space unlike [5].
Organization:
This paper is organized as follows.
Section 2 summarizes basic notions around the optimal transport theory.
Section 3 introduces the IV multivariate quantile regression model.
Section 4 formally states the local identification result for the model.
In Section 5, we provide a broadly applicable identification theorem that is used in Section 6 to show the identification result for our model.
Section 7 discusses related issues.
Most proofs appear in Appendix A.
3 IV Multivariate Quantile Model
Let be a -dimensional potential outcome vector, an endogenous variable (treatment), an exogenous covariate vector, and an IV.
Let be a compact convex subset with a piecewise boundary
Let be a reference probability measure of which support is
Assume is absolutely continuous with respect to the -dimensional Lebesgue measure.
Our model extends [7] so that it accounts for multidimensional outcome vectors.
The following is the primitives.
Assumption 3.1.
The random variables satisfy the following conditions with probability one:
1.
For each it holds that and there exists a function such that is continuously differentiable in the first argument and its derivative is symmetric and positive definite on
2.
For each
3.
for some unknown function and random vector
4.
Conditional on are identically distributed.
5.
The observed random vector consists of and
The difference from [7] appears in (1).
Recall from Section 2 that the existence of the multivariate quantile function is not restrictive.
Essentially, (1) assumes the smoothness of and the strict positive definiteness of its derivative.
Also, the continuity of imposed in (1) implies that the range of is compact in and so is the support of conditional on
This restriction can be relaxed, but we keep it for simplicity.
Condition (2) implies that the IV is independent of potential outcomes conditional on covariates.
Condition (3) is a weak restriction that allows a broad class of assignment rules of treatments.
In particular, in (3) can be correlated with potential outcomes.
Condition (4) is called the rank similarity, which is usually assumed in the literature.
This is weaker than the rank invariance assumption that does not change with assigned treatments.
For more details of the assumptions, see [7] and [9].
For the rank similarity condition (4), see also [21] and references therein.
Notice that by condition (1), the observed outcome is written as where
The following representation is the main testable implication of the model.
Theorem 3.1.
Suppose Assumption 3.1 holds.
Then, we have with probability one, it holds that for each measurable
where
As the LHS of (4) is identified as a function of the equation gives an identification restriction.
That is, any candidate of structural functions necessarily satisfies (4) in which is replaced by
In the latter section, we consider conditions under which is the only functions that satisfy the equation.
For this restriction is boiled down to a much simpler form.
Consider the case of and
Since is strictly increasing and the class forms a -system that generates the Borel -algebra on the equation (4) is equivalent to
for each where the second equality is considered in Theorem 1 of [7].
Notice that conditional on the LHS is a function of numbers for each
This is a critical simplification that occurs only in the one-dimensional case in the sense that the set which is intrinsically an infinite-dimensional object, is pined down by the scalar
Thanks to this dimension reduction, in order to consider the identification of structural functions, one can fix a certain quantile level and focus on a finite-dimensional system of equations, as [7] do.
In higher dimensions, however, it is practically impossible to find a -system with such a nice property.
For example, for and consider a rectangle where
The image of the structural function is not determined only by even under the cyclical monotonicity assumption, because the image set is nonlinear in general.
To characterize it, one need to know the set of the values of at all the boundary points of which is clearly an infinite-dimensional object.
This fact makes the problem for completely distinct and harder to solve than the one-dimensional case.
The fact that the image of the structural function is not pined down by a single point in the multidimensional setting implies that just considering the identification of the structural function at a certain quantile level makes no sense unlike the case of
We consider the identification in an infinite-dimensional function space in the next section.
4 Local Identification of Structural Functions
In this section, we consider the identification problem of the structural functions under Assumption 3.1.
For the notation simplicity, let by suppressing and all the following analysis should be understood as conditional on
Notice that this treatment does not affect identification results because the distributions of observables conditional on are identifiable.
The representation (4) of Theorem 3.1 implies by the change-of-variables formula that the true structural functions solve the following equation
for where
For a scalar function a vector-valued function and a matrix-valued function we define three types of supremum norms as follows:
For a compact set in a Euclidean space, we denote the set of continuous functions on by
Also, let be the set of functions on that are -times continuously differentiable on and the derivative can be continuously extended to
Our identification result shows that is the locally unique solution to (5) in a certain function class.
To state the result formally, let
Then, is a normed space.
Also, let be the set of vectors of functions satisfying (1) of Assumption 3.1, i.e.,
Fix and
As the parameter space, we consider
where is the largest eigenvalue of and define similarly.
Since we are interested in identification, we assume the correct specification.
Assumption 4.1.
Since for has a symmetric positive definite derivative, by Lemma 2.1, there is a convex function on such that and is unique up to an additive constant.
Condition (ii) on the eigenvalues that appears in the definition of controls the convexity and smoothness of
To see this, recall that a smooth convex function is called -strongly convex () if is positive semi-definte, and that it is called -strongly smooth () if is so.
Geometrically, the strong convexity requires that should not be too flat while the strong smoothness prevents it from being too curved.
In terms of the first derivative these conditions make neither too gentle nor too steep.
Also, if satisfies then it is -strongly convex and -strongly smooth.
Given these, we can intuitively understand that Assumption 4.1 assumes that is moderately steep.
Condition (iii) in the definition of excludes pathological candidates of structural functions.
This type of condition is often necessary for nonparametric identification.
Section 2.3 of [5] discusses an example in which identification fails if an analogous condition to (iii) is removed.
We further impose three regularity conditions.
Assumption 4.2.
For each the support of is a convex bounded set with a piecewise boundary
Assumption 4.3.
For each
Assumption 4.4.
For each the support of is identified, that is,
Assumption 4.2 requires that the support of a potential outcome should be regular.
The convexity of the support is necessary for many purposes in the optimal transport theory (see, for example, [29] and [16]).
The smoothness of the boundary is rarely a problem in practice, as it allows for kinks at some points.
Combined with Assumption 4.2, Assumption 4.3 implies that is bounded.
The identifiability of the support of in Assumption 4.4 puts a nontrivial restriction on structural functions.
Recall that each of is the optimal transport map from to
Unlike the common setup, we do not know the target distribution due to the endogeneity, so that there is not much to be gained from the standard optimal transport theory.
Nevertheless, knowing the support of the target distribution rules out many candidates of structural functions, which we exploit for identification.
We discuss further details on this in Section 6.
In what follows in this section, we focus on the case of to develop the idea clearly.
A more general result is given in Appendix B.
The following is a key assumption.
Assumption 4.5.
For
holds.
The following is our main theorem.
Theorem 4.1.
Suppose that Assumptions 3.1, 4.1, 4.2, 4.3, 4.4, 4.5 hold.
Then, the structural functions are locally identified in i.e., there exists such that if with solves the system (5), then holds.
Assumption 4.5 is a novel assumption that we cannot find in the literature.
Roughly speaking, this condition requires () to be an informative IV for ().
For instance, if is a perfect instrument of that is, then the condition (6) holds since
As another example, consider the case where is uniformly distributed independently of
In this case, the condition (6) is reduced to
which holds if and are large enough.
Note that Assumption 4.5 requires a positive correlation between and so the labeling of the IV should be switched in favor of the assumption.
When [7] show that the identification holds if the matrix
is full rank.
Assumption 4.5 is stronger than this condition.
Too see this, notice
Thus, equation (6) implies
and therefore, the determinant of the matrix (9) is strictly positive.
In particular, the matrix is full rank.
The difference between [7] and us is that our proof relies on the positive definiteness of a certain matrix that depends on (see Lemma 6.3 below) while just the full rankness is required in [7].
As their proof idea works only when the fact that the full rank condition is replaced by Assumption 4.5 can be seen as the cost of taking multidimensional potential outcomes into account.
Of course, our proof can be easily modified to show Theorem 4.1 for under the full rank condition of [7], which is discussed in Section 7.
Assumption 4.5 corresponds to the condition known as the ellipticity or the coercivity in the elliptic PDE literature.
It is almost always assumed to discuss the well-posedness or the regularity of solutions to elliptic PDEs.
See [18] for details.
As the dimension increases, Assumption 4.5 becomes more demanding.
This is quite natural, as one needs to identify higher dimensional structural functions.
However, it is surprising that the minimum size of the support of the IV, needed for the identification is independent of the dimension of the potential outcomes, as long as the IV is informative enough.
5 Local Identification for Systems of Equations
The proof of Theorem 4.1 relies on a more general identification result (Theorem 5.1), which is of independent interest.
We devote this section to the identification result that can be applied to parameters that are characterized by the solution of a system of equations.
Let be normed spaces, be any subset, and be a finite set.
For fix and
We consider the following system of equations
Suppose is a solution to (10).
We are interested in whether is the locally unique solution.
To formulate deviations from we define the tangent set at as
The set specifies deviations in that can be taken from
Notice that if then
We also define the set of normalized tangent vectors as
Remember that when and are finite-dimensional, the locally unique solvability of the system of equations is given by the implicit function theorem, which assumes the differentiability and the full rankness of the Jacobian matrix.
As our identification result is an analogue statement for infinite-dimensional parameters, we impose assumptions corresponding to these two.
The first assumption is the differentiability of at
Assumption 5.1.
For there exists such that
where if the set over which the supremum is taken is empty, the value reads zero.
It is clear that if Assumption 5.1 is reduced to the standard Fréchet differentiability.
For strict subsets of it is weaker than the Fréchet differentiability.
Note that may not be a vector space, but Assumption 5.1 works even in this case.
Unlike the Fréchet differentiability, the linearity of is not required.
The next assumption is an infinite-dimensional version of the full rankness of the Jacobian matrix.
Assumption 5.2.
It holds that
The following lemma provides an equivalent representation of Assumption 5.2.
Lemma 5.1.
Let
Assumption 5.2 holds if and only if there exists such that for all
Now, we are ready to state the main theorem in this section.
Theorem 5.1.
Under Assumptions 5.1 and 5.2, is the locally unique solution to (10).
That is, there exists such that for all if then is not a solution to the system.
This result is closely related to Section 2 of [5].
Indeed, Their Theorems 1 can be shown from Theorem 5.1.
See Appendix C for details.
Our theorem has a different focus than their result in that it covers the local identification problem for systems of equations and that the domain of functionals is not necessarily be a Banach space.
6 Proof of Theorem 4.1
In this section, we prove Theorem 4.1.
To do so, we apply Theorem 5.1.
Recall that is a normed space and that
Let be the space of finite signed Borel measures on equipped with the total variation norm:
Then, is a normed space.
For define as
Then, the identification restriction (5) is written as for all
We apply Theorem 5.1 to this system of equations where and
The normalized tangent set that appears in Section 5 is
We need to check Assumptions 5.1 and 5.2.
Remember that for a smooth vector field the divergence of at is defined as
The following lemma states that is differentiable in the sense of Assumption 5.1.
Lemma 6.1.
Under Assumptions 4.1 and 4.3, the operator satisfies Assumption 5.1 with
for
Next, we show the Assumption 5.2 by contradiction.
Suppose there is a sequence such that
as
By definition, there exists such that
As and are optimal transport maps, they are written as and for some convex functions and
By adding a constant, we may assume without loss of generality.
Then satisfies and
Since there exists such that it holds that for any
where
Hence, we have
where the last inequality holds because
For the simplicity, assume below.
Appendix B discusses the case of general treatments and IVs.
Let
As we have which is independent of
Also, it holds that and that
By considering a function in the definition (13) of the total variation norm, the integration by parts implies
where
where is the outward normal unit vector of at and is the -dimensional surface measure on
The following lemma, which exploits Assumption 4.4, the identifiability of the support of is a key step in our proof.
Remember that since is continuous, implies
Lemma 6.2.
Suppose that Assumption 4.4 holds.
For and define
Then, at any point is an outward normal vector of
By Lemma 6.2, we have
Since the support is convex by Assumption 4.2, it holds that for which implies
Thus, we have and therefore,
Taking the sum over yields
Now, we restrict our attention to the case of
The following lemma shows the positive definiteness of the integrand of the RHS of (15).
Lemma 6.3.
Suppose
Assumption 4.5 implies that for
holds for
By Lemma 6.3, the integrand in the RHS of (15) is nonngetive under Assumption 4.5.
Since the LHS of (15) converges to zero by (14) and we have
for almost all by taking a subsequence if necessary.
Recall that and since
Then, the sequence is uniformly bounded and uniformly equicontinuous, and therefore, there exists a uniformly converging subsequence by the Arzelà–Ascoli theorem.
Let be the limit function.
By equation (16), almost everywhere, and it holds everywhere, as is continuous.
Hence, the uniform convergence implies up to a subsequence.
However, this contradicts to the fact that which particularly implies
Thus, Assumption 5.2 is satisfied.
Now, Theorem 4.1 is an immediate consequence of Theorem 5.1.